OpenStudy (anonymous):

how to factor a polynomial with a fraction. EX. x^2+7x+11

OpenStudy (anonymous):

well , you can do this x(x+7) +11

OpenStudy (anonymous):

I think your best bet is to complete the square to solve for vertex or use the quadratic formula because its not factorable

OpenStudy (anonymous):

If we have to use a perfect square, how would we do that?

OpenStudy (anonymous):

You set it up like this: (x^2 + 7x) + 11 You divide the coefficient of x which is 7 by 2. Square (7/2) and add it to both sides

OpenStudy (anonymous):

So now you have (x^2 + 7x + 49/4) -49/4 + 44/4

OpenStudy (anonymous):

(x + 7/2)^2 -5/4

OpenStudy (anonymous):

Wait but how do you simplify x^2+7x+49/4, because I got up to this point, I dont' get what to do after...

OpenStudy (anonymous):

You have to consider what multiply to 49/4 and add up to 7

OpenStudy (anonymous):

okay because its an equation whatever you do to the right side must be applied to the left side...if initially you set up your equation equal to 0 then when you added 49/4 you will also add 49/4 to the right side

OpenStudy (anonymous):

x^2 + 7x + 49/4 + 11 = 49/4

OpenStudy (anonymous):

(x^2 + 7x + 49/4) becomes (x + 7/2)^2

OpenStudy (anonymous):

And then you subtract 11 which is 44/4 from 49/4 = 5/4

OpenStudy (anonymous):

So you have (x + 7/2)^2 - 5/4 = 0...do you understand it?

OpenStudy (anonymous):

Oh! So the middle number is what you add to get and the last part of the equation (49/4) is what you multiply to get.

OpenStudy (anonymous):

Yes, that's correct

OpenStudy (anonymous):

Thank you you have been very helpful! is there a way I can +1 you or something? I just signed up so I have no clue how this works :S

OpenStudy (anonymous):

you can "Good Answer" me

OpenStudy (anonymous):

yes that's correct! Glad you understand